OpenStudy (anonymous):

how can I find all complex solutions of (4/x^2-3x)-(1/x^2-9)=0

hero (hero):

\[\frac{4}{x^2 - 3x} - \frac{1}{x^2 - 9} = 0\]

hero (hero):

\[\frac{4}{x^2 - 3x} = \frac{1}{x^2 - 9}\]

hero (hero):

\[4(x^2 - 9) = 1(x^2 - 3x)\] \[4x^2 - 36 = x^2 - 3x\] \[3x^2 + 3x - 36 = 0\] \[x^2 + x - 12 = 0\] \[(x + 4)(x - 3) = 0\] \[x = {-4,3}\] I don't think this has any complex solutions.

hero (hero):

@Diana0920

OpenStudy (anonymous):

@Hero thanks (: